John Bassett, the owner of the Tampa Bay Bandits, offered a deal where if they sell out the stadium during the 1985 season one fan will win a million dollar annuity. The catch, they will be paid $50,000 a year for 20 years starting in 20 years. This is a true story!

Let’s say you are John Bassett and you go to the bank. The bank offers you an APR for this payout annuity of 4.2% compounded annually.

To soften the blow the bank allows you to make monthly payments towards the annuity for the next 20 years at 3.8% compounded monthly.

A) Calculate the needed Nest Egg in 20 years when the annuity begins making annual payments. What is the formula?_________

B) Calculate the needed monthly payments for the next 20 years to reach this nest egg.________
Answer #1

A) Calculate the needed Nest Egg in 20 years when the annuity begins making annual payments. What is the formula?_________
=50000/(4.2%/12)*(1-1/(1+4.2%/12)^(12*20))=8109369.636

B) Calculate the needed monthly payments for the next 20 years to reach this nest egg.________
=8109369.636*(3.8%/12)/((1+3.8%/12)^(12*20)-1)=22611.11355